# Fast algorithms to improve fair information access in networks

### v1

@@ Appendix B Pseudocode for Algorithms
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:Select $a_{1},a_{2},a_{3},a_{4},a_{5}$ to maximize number of networks in the corpus for which performance of some $a_{i}$ is $\geq 80\%$ of Myopic.
6:$D\leftarrow\{a_{1},a_{2},a_{3},a_{4},a_{5}\}$
7:// ask the oracle $O$ which algorithm to use.
8:$a_{best}\leftarrow O(G,D)$
9:// Run the algorithm to get $S$.
10:$S=S\cup a_{best}(G,s)$
11:Return $S$.
List of Algorithms 11 Fast Ensemble (Oracle)
@@
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:Select $a_{1},a_{2},a_{3},a_{4},a_{5}$ to maximize number of networks in the corpus for which performance of some $a_{i}$ is $\geq 80\%$ of Myopic.
6:$D\leftarrow\{a_{1},a_{2},a_{3},a_{4},a_{5}\}$
7:Train Random Forest Classifier $M$ to select the best algorithm from $D$ based on network topology, using data from the network corpus.
8:// Ask $M$ which algorithm to use.
9:$a_{best}\leftarrow M(G,D)$
10:Run the algorithm to get $S$
11:$S=S\cup a_{best}(G,s)$
12:Return $S$.
List of Algorithms 12 Meta-learner
@@ Appendix C Supplementary Results
Figure S2: Mean performance of the intervention algorithms on each domain in the corpus, under (a) low spreadability and (b) high spreadability settings. Medium spreadability results are given in Fig. 5.
@@
Figure S3: Performance of intervention algorithms on the network corpus, relative to Myopic and sorted in ascending order by network size within each domain. (a): low spreadability, $12\%$ of networks have no algorithm better than or within $80\%$ of Myopic’s performance; (b): high spreadability, showing very similar results to medium spreadability (Fig. 6). “Equivalent” defined as within one standard error of $\beta$ for Myopic; typically about 0.001.
Figures S4 and S5 tabulate the counts of how often a particular algorithm was the best performing (highest $\beta$) and plot them as a function of network mean degree $\langle k\rangle$, for low and high spreadability, respectively. Here we see clearly that Myopic is by far the best algorithm in high spreadability (similar to results for medium, shown in Fig. 7), but is much less so under low spreadability. We believe this difference is attributable to the imprecision of ProbEst in the low spreadability setting.
Figure S4: Best-performing algorithm on a network vs average degree of the network under low spreadability. Each network’s corresponding single most-performant algorithm is plotted with a circle. Total number of circles per line is given on the right-hand side, indicating how many times an algorithm was the best-performing one across the entire corpus.
Figure S5: Best-performing algorithm on a network vs average degree of the network under high spreadability. Each network’s corresponding single most-performant algorithm is plotted with a circle. Total number of circles per line is given on the right-hand side, indicating how many times an algorithm was the best-performing one across the entire corpus.
@@
This multi-class classification task is challenging. Fine-tuned models come out with an accuracy of $\approx 0.51$. Figs. S6 and S7 show the learned feature importances and confusion matrix for the medium spreadability setting (similar results for low and high spreadability models). However, generally, incorrect algorithm choices by the model tend to select an alternative that is almost as good as the true best, in terms of performance $\beta$, and so classification errors often do not translate into large losses in performance.
@@
From the classification task, we find that the domain of a network contributes the least to the prediction accuracy, followed by number of nodes $n$. Other network features are roughly equally important (Fig. S6). This results implies that network domain does not contribute much marginal information beyond a network’s features, as has been found previously by Ref. ikehara2017 . That is, network features encode much the same information as network domain.
Figure S6: Meta-learner random forest feature importance for predicting which of five algorithms is the best choice for a given network. Higher values mean a feature contributes more to the predicted performance value.
Figure S7: Confusion matrix of the best algorithm prediction model for the meta-learner.
@@
The accuracy of the meta-learning model under medium and high spreadability (Figs. 9 and S8) are similar, but with slightly larger performance loss and substantially larger runtime speedup in the case of high, than in medium.
@@
Unsurprisingly, the fast ensemble algorithms (which use an oracle to choose the best heuristic for a given network) in both medium and high spreadability settings outperform the meta-learner. However, the gap between fast ensemble and meta-learning is only about 10-20%, indicating that while there is room for improvement in the meta-learner’s classifier, the optimal performance is not far off. Moreover, the fast ensemble does outperform Myopic on more cases than does the meta-learning, but again, the gap is not enormous.
@@
Figure S8: Performance difference vs. speedup for the meta-learning algorithm relative to Myopic under high spreadability, with marginal histograms. Results averaged over $1000$ meta-learner runs, and extreme outliers have been removed for visualization purposes only. Average performance loss relative to Myopic is $28.20\pm 25.80$, with an average speedup factor of $133.35\pm 79.32$, but for 22 networks (12.9%), the meta-learner outperforms Myopic.
@@
Figure S9: Performance difference vs. speedup for the fast ensemble algorithm relative to Myopic under medium spreadability, with marginal histograms. Extreme outliers have been removed for visualization purposes only. Average performance loss relative to Myopic is $9.34\pm 28.34\%$, with an average speedup factor of $70.66\pm 85.65$, but for 48 networks (28.1%), the fast ensemble outperforms Myopic.
@@
Figure S10: Performance difference vs. speedup for the fast ensemble algorithm relative to Myopic under high spreadability, with marginal histograms. Extreme outliers have been removed for visualization purposes only. Average performance loss relative to Myopic is $11.11\pm 21.8\%$, with an average speedup of $101.88\pm 101.27$, but for 47 networks (27.6%), the fast ensemble algorithm outperforms Myopic.

### v2

@@ 6       $S=S\cup\{y\}$
Input : $G=(V,E),s$.
Output : $S\subseteq V$ with $|S|=s+1$.
1
2Initialize $S$ with a random node.
3 $S=\{s_{0}\in_{R}V\}$
4Select $a_{1},a_{2},a_{3},a_{4},a_{5}$ to maximize number of networks in the corpus for which performance of some $a_{i}$ is $\geq 80\%$ of Myopic
5 $D\leftarrow\{a_{1},a_{2},a_{3},a_{4},a_{5}\}$
// Ask the oracle $O$ which algorithm to use
6 $a_{\rm best}\leftarrow O(G,D)$
// Run the algorithm to get $S$
7 $S=S\cup a_{\rm best}(G,s)$
return $S$
Algorithm 11 Fast Ensemble (Oracle)
@@ 7 $S=S\cup a_{\rm best}(G,s)$
Input : $G=(V,E),s$.
Output : $S\subseteq V$ with $|S|=s+1$.
1
2Initialize $S$ with a random node.
3 $S=\{s_{0}\in_{R}V\}$
4Select $a_{1},a_{2},a_{3},a_{4},a_{5}$ to maximize number of networks in the corpus for which performance of some $a_{i}$ is $\geq 80\%$ of Myopic
5 $D\leftarrow\{a_{1},a_{2},a_{3},a_{4},a_{5}\}$
6Train Random Forest Classifier $M$ to select the best algorithm from $D$ based on network topology, using data from the network corpus
7
// Ask $M$ which algorithm to use
8 $a_{\rm best}\leftarrow M(G,D)$
// Run the algorithm to get $S$
9 $S=S\cup a_{\rm best}(G,s)$
return $S$
Algorithm 12 Meta-learner
Figure S1. Mean performance of the intervention algorithms on each domain in the corpus, under (a) low spreadability and (b) high spreadability settings. Medium spreadability results are given in Figure 6.
Figure S2. Performance of intervention algorithms on the network corpus, relative to Myopic and sorted in ascending order by network size within each domain. (a): low spreadability, $12\%$ of networks have no algorithm better than or within $80\%$ of Myopic’s performance; (b): high spreadability, showing very similar results to medium spreadability (Figure 7). “Equivalent” defined as within one standard error of $\beta$ for Myopic; typically about 0.001.
@@ Appendix C Supplementary Results
Figures S3 and S4 tabulate the counts of how often a particular algorithm was the best performing (highest $\beta$) and plot them as a function of network mean degree $\langle k\rangle$, for low and high spreadability, respectively. Here we see clearly that Myopic is by far the best algorithm in high spreadability (similar to results for medium, shown in Figure 8), but is much less so under low spreadability. We believe this difference is attributable to the imprecision of ProbEst in the low spreadability setting.
Figure S3. Best-performing algorithm on a network vs average degree of the network under low spreadability. Each network’s corresponding single most-performant algorithm is plotted with a circle. Total number of circles per line is given on the right-hand side, indicating how many times an algorithm was the best-performing one across the entire corpus.
Figure S4. Best-performing algorithm on a network vs average degree of the network under high spreadability. Each network’s corresponding single most-performant algorithm is plotted with a circle. Total number of circles per line is given on the right-hand side, indicating how many times an algorithm was the best-performing one across the entire corpus.
Figure S5. Meta-learner random forest feature importance for predicting which of five algorithms is the best choice for a given network. Higher values mean a feature contributes more to the predicted performance value.
Figure S6. Confusion matrix of the best algorithm prediction model for the meta-learner.
@@
This multi-class classification task is challenging. Fine-tuned models come out with an accuracy of $\approx 0.51$. Figs. S5 and S6 show the learned feature importances and confusion matrix for the medium spreadability setting (similar results for low and high spreadability models). However, generally, incorrect algorithm choices by the model tend to select an alternative that is almost as good as the true best, in terms of performance $\beta$, and so classification errors often do not translate into large losses in performance.
@@
From the classification task, we find that the domain of a network contributes the least to the prediction accuracy, followed by number of nodes $n$. Other network features are roughly equally important (Figure S5). This result implies that network domain does not contribute much marginal information beyond a network’s features, as has been found previously by Ikehara and Clauset [28]. That is, network features encode much the same information as network domain.
@@
The accuracy of the meta-learning model under medium and high spreadability (Figs. 10 and S7) are similar, but with slightly larger performance loss and substantially larger runtime speedup in the case of high, than in medium.
@@
Unsurprisingly, the fast ensemble algorithms (which use an oracle to choose the best heuristic for a given network) in both medium and high spreadability settings outperform the meta-learner. However, the gap between fast ensemble and meta-learning is only about 10-20%, indicating that while there is room for improvement in the meta-learner’s classifier, the optimal performance is not far off. Moreover, the fast ensemble does outperform Myopic on more cases than does the meta-learning, but again, the gap is not enormous. We note that timing results for some algorithms (both those using ProbEst and the BFS-based class of new algorithms; see Section 3) may be impacted by our use of an adjacency matrix (rather than adjacency list) representation for all networks.
@@
Figure S7. Performance difference vs. speedup for the meta-learning algorithm relative to Myopic under high spreadability, with marginal histograms. Results averaged over $1000$ meta-learner runs, and extreme outliers have been removed for visualization purposes only. Average performance loss relative to Myopic is $28.20\pm 25.80$, with an average speedup factor of $133.35\pm 79.32$, but for 22 networks (12.9%), the meta-learner outperforms Myopic.
@@
Figure S8. Performance difference vs. speedup for the fast ensemble algorithm relative to Myopic under medium spreadability, with marginal histograms. Extreme outliers have been removed for visualization purposes only. Average performance loss relative to Myopic is $9.34\pm 28.34\%$, with an average speedup factor of $70.66\pm 85.65$, but for 48 networks (28.1%), the fast ensemble outperforms Myopic.
@@
Figure S9. Performance difference vs. speedup for the fast ensemble algorithm relative to Myopic under high spreadability, with marginal histograms. Extreme outliers have been removed for visualization purposes only. Average performance loss relative to Myopic is $11.11\pm 21.8\%$, with an average speedup of $101.88\pm 101.27$, but for 47 networks (27.6%), the fast ensemble algorithm outperforms Myopic.
Appendix D Results on Larger Networks
$n$ $m$ $k$ $C$ $l_{max}$ $\sigma^{2}$
Email (EU) 34,203 96,669 15.87 0.04 7 770
Google+ (2013) 86,785 688,602 5.65 0.27 47 1,809
Table S2. Statistics of Email (EU) and Google+ (2013) showing number of nodes $n$, number of edges $m$, average node degree $k$, clustering coefficient $C$, diameter $l_{max}$, and the variance $\sigma^{2}$ of the degree distribution.
In Table S2, we present summary statistics for the two larger networks used in Section 4.4. In Figure S10, we evaluate the performance of the five algorithms from $\mathcal{X}_{\text{high}}$ as well as that of Myopic. When evaluating the performance of these algorithms, it is important to note that due to the size of the networks, a large increase in $\pi_{\text{min}}$ is not expected with any algorithm when choosing only 10 seeds. Additionally, when looking at the minimum access probability as seeds are added, there are small decreases in $\pi_{\text{min}}$ resulting from the imprecise nature of ProbEst. This is a limitation on our ability to evaluate these algorithms, but it is not a restriction on the algorithms themselves. That being noted, three out of the five algorithms (MinDegree_hc, Gonzalez and LeastCentral) produce $\beta$ values close to, and in some cases even greater than, Myopic (Figure S10). Although the $\beta$ values are predictably low for all algorithms, including Myopic, there is a general positive trend with most algorithms. From these results, we hypothesize that when scaled to large networks, the speedup of the algorithms is sustained, while seed choice quality remains comparable to Myopic for at least some algorithms.
Figure S10. Minimum access probability vs. seed set size under the six algorithms in $\mathcal{X}_{\rm high}$, evaluated on the two larger networks, Email (EU) and Google+ (2013) under high spreadability. Small decreases in $\pi_{\text{min}}$ are discussed in Appendix D. The panel on the right shows $\beta$, the slope of the line of best fit, for each algorithm.